# Local-global principle for triangularizability and diagonalizability of matrices

### v1

@@ 5. The stratified Brauer–Manin obstruction
Lemma 5.15.
[…]
Then each $V_{\textbf{r}}$ is a closed subscheme of $X_{M}$. Moreover, we have
$X_{M}=\bigcup_{\textbf{r}\in\mathfrak{R}_{m,n}}V_{\textbf{r}}.$
Proof.
For each $1\leq i\leq n-1$, define
$r_{i}:=\min\left\{t\mid\mathrm{rank}(A^{m+i-1,t})<t\right\}.$
[…]
Therefore, we have
Proof of Theorem 5.4.
According to Lemma 5.14 and Lemma 5.15, we only need to prove that for any $\mathbf{r}\in\mathfrak{R}_{m,n}$, $V_{\textbf{r}}$ is a smooth and $k$-rational variety satisfying $\overline{V_{\textbf{r}}(k)}=V_{\textbf{r}}(\textbf{A}_{k})_{\bullet}^{\textup{Br}}$.
Our proof proceeds by induction on $n$. Assume that $\mathbf{r}=(r_{l})_{1\leq l\leq n}\in\mathfrak{R}_{m,n}$ (set $r_{0}=0$ when $n=2$). In fact, when $r_{n-1}<m+n-1$, one can easily check that $V_{\textbf{r}}$ is just $W^{\prime}_{\textbf{s}}$ in Lemma 5.12 where $\textbf{s}=(r_{1},...,r_{n-1},r_{n-1})\in\mathfrak{S}_{m,n}$ is a well-defined element, so $V_{\textbf{r}}$ is a smooth and $k$-rational variety satisfying $\overline{V_{\textbf{r}}(k)}=V_{\textbf{r}}(\textbf{A}_{k})_{\bullet}^{\textup{Br}}$. Hence we only need to consider the case where $r_{n-1}=m+n-1$.
$\det(T_{i,j})=T_{m+n,m+n}\cdot C_{m+n,m+n},$
where the cofactor $C_{m+n,m+n}$ is of the form
So one can easily check that when $n=2$, we have $V_{\textbf{r}}\cong\mathrm{GL}_{m+1}\times\mathbb{G}_{m}\times\mathbb{A}^{m+1}_{k}$, hence is a smooth and $k$-rational variety satisfying $\overline{V_{\textbf{r}}(k)}=V_{\textbf{r}}(\textbf{A}_{k})_{\bullet}^{\textup{Br}}$ by Lemma 2.2; when $n\geq 3$, we have $V_{\textbf{r}}\cong V_{\textbf{r}^{\prime}}\times\mathbb{G}_{m}\times\mathbb{A}^{m+n-1}_{k}$, where $\textbf{r}^{\prime}=(r_{1},r_{2},...,r_{n-2})\in\mathfrak{R}_{m,n-1}$ is a well-defined element, hence is a smooth and $k$-rational variety satisfying $\overline{V_{\textbf{r}}(k)}=V_{\textbf{r}}(\textbf{A}_{k})_{\bullet}^{\textup{Br}}$ by the induction hypothesis and Lemma 2.1. This completes the proof. ∎
Remark 5.16.
Notably, the authors have also established that Conjectures 5.2 and 5.3 hold for all $M$ with order $\leq 5$. All the results so far are obtained using the fibration method. However, different fibrations are employed for different cases, making it difficult to unify them. Consequently, the validity of these conjectures in full generality remains open.

### v2

@@ 5. The stratified Brauer–Manin obstruction
For each $1\leq i\leq n-1$, define
$r_{i}:=\min\left\{t\mid\mathrm{rank}(A^{m+i-1,t})<t\right\}.$
[…]
Therefore, we have
Proof of Theorem 5.4.
According to Lemma 5.14 and Lemma 5.15, we only need to prove that for any $\mathbf{r}\in\mathfrak{R}_{m,n}$, $V_{\textbf{r}}$ is a smooth and $k$-rational variety satisfying $\overline{V_{\textbf{r}}(k)}=V_{\textbf{r}}(\mathbf{A}_{k})_{\bullet}^{\textup{Br}}$.
Our proof proceeds by induction on $n$. Assume that $\mathbf{r}=(r_{l})_{1\leq l\leq n}\in\mathfrak{R}_{m,n}$ (set $r_{0}=0$ when $n=2$). In fact, when $r_{n-1}<m+n-1$, one can easily check that $V_{\textbf{r}}$ is just $W^{\prime}_{\textbf{s}}$ in Lemma 5.12 where $\textbf{s}=(r_{1},...,r_{n-1},r_{n-1})\in\mathfrak{S}_{m,n}$ is a well-defined element, so $V_{\textbf{r}}$ is a smooth and $k$-rational variety satisfying $\overline{V_{\textbf{r}}(k)}=V_{\textbf{r}}(\mathbf{A}_{k})_{\bullet}^{\textup{Br}}$. Hence we only need to consider the case where $r_{n-1}=m+n-1$.
$\det(T_{i,j})=T_{m+n,m+n}\cdot C_{m+n,m+n},$
where the cofactor $C_{m+n,m+n}$ is of the form
So one can easily check that when $n=2$, we have $V_{\textbf{r}}\cong\mathrm{GL}_{m+1}\times\mathbb{G}_{m}\times\mathbb{A}^{m+1}_{k}$, hence is a smooth and $k$-rational variety satisfying $\overline{V_{\textbf{r}}(k)}=V_{\textbf{r}}(\mathbf{A}_{k})_{\bullet}^{\textup{Br}}$ by Lemma 2.2; when $n\geq 3$, we have $V_{\textbf{r}}\cong V_{\textbf{r}^{\prime}}\times\mathbb{G}_{m}\times\mathbb{A}^{m+n-1}_{k}$, where $\textbf{r}^{\prime}=(r_{1},r_{2},...,r_{n-2})\in\mathfrak{R}_{m,n-1}$ is a well-defined element, hence is a smooth and $k$-rational variety satisfying $\overline{V_{\textbf{r}}(k)}=V_{\textbf{r}}(\mathbf{A}_{k})_{\bullet}^{\textup{Br}}$ by the induction hypothesis and Lemma 2.1. This completes the proof. ∎
Remark 5.16.
Notably, the authors have also established that Conjecture 5.2 and Conjecture 5.3 hold for all $M$ with order $\leq 5$. All the results so far are obtained using the fibration method. However, different fibrations are employed for different cases, making it difficult to unify them. Consequently, the validity of these conjectures in full generality remains open.
Appendix A Proof of Lemma 5.14
In this appendix, we provide the detailed proof of Lemma 5.14.
Proof of Lemma 5.14.